# Anytime-valid, Bayes-assisted, Prediction-Powered Inference

### v1

@@ 7 Discussion
$\overline{Y}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+\varepsilon_{t}\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ where\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\varepsilon_{t}=o\left(\frac{1}{t^{1-1/(2+\delta)}}\right)=o\left(\frac{1}{\sqrt{t\log t}}\right).$
The result follows. ∎
S4.3 Tighter strong coupling for control-variate estimators
Here, we state tighter equivalents of Proposition 4.1 and Proposition 4.2. We consider the same notations as in Section 4.1, with the additional assumption that $\mathbb{E}|U|^{2+\delta},\mathbb{E}|V|^{2+\delta}<\infty$ for some $\delta>0$.
Proposition S4 (Tighter version of Proposition 4.1).
Assume $\mathbb{E}|U|^{2+\delta}$ and $\mathbb{E}|V|^{2+\delta}<\infty$ for some $01>\delta>$ 0. We have, almost surely as $n\to\infty$,
@@
$\widehat{\gamma}^{\mathrm{cv+}}$ $=\widehat{\gamma}^{\mathrm{cv}}_{\lambda^{*}}+o\left(\frac{1}{\sqrt{n\log n}}\right)=\overline{V}-\lambda^{*}(\overline{U}-\widehat{\mu})+o\left(\frac{1}{\sqrt{n\log n}}\right).$ (S42)
Proof.Let
$\epsilon_{n}$ $=\widehat{\gamma}^{\mathrm{cv+}}-\widehat{\gamma}_{\lambda^{*}}^{\mathrm{cv}}$
$=\widehat{\gamma}^{\mathrm{cv+}}-(\overline{V}-\lambda^{*}(\overline{U}-\widehat{\mu}))$
$=(\lambda^{*}-\widehat{\lambda})(\overline{U}-\mu)-(\lambda^{*}-\widehat{\lambda})(\widehat{\mu}-\mu).$
@@
$|\epsilon_{n}|\leq|\lambda^{*}-\widehat{\lambda}|(|\overline{U}-\mu|+|\widehat{\mu}-\mu|).$
By the Marcinkiewicz-Zygmund strong law of large numbers (Theorem S2) with $p=1+\delta/2$, we have that
$|\widehat{\lambda}-\lambda^{*}|=o\left(n^{-1/(1+\delta/2)}\right)=o\left(\frac{1}{n^{2/(2+\delta)}}\right)\text{ a.s. as }n\to\infty$
@@
$|\overline{U}-\mu|=O\left(\sqrt{\frac{\log\log n}{n}}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ |\widehat{\mu}-\mu|=O\left(\sqrt{\frac{\log\log n}{n}}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ a.s. as }n\to\infty.$ (S43)
@@
$|\epsilon_{n}|=O\left(\frac{1}{n^{2/(2+\delta)}}\sqrt{\frac{\log\log n}{n}}\right)=o\left(\frac{1}{\sqrt{n\log n}}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ a.s. as }n\to\infty.$ (S44)
Proposition S5 (Tighter version of Proposition 4.2, KMT for CVEs).
Assume that $|\frac{n}{N_{n}}-r|=O(1/n^{1-a})$ with $0<a<2/(2+\delta)$ for some $r\in[0,1]$. There exist i.i.d. Gaussian random variables $(W^{\mathrm{cv}}_{i})_{i\geq 1}$ with mean $\gamma$ and variance
such that, almost surely as $n\to\infty$,
$\widehat{\gamma}_{\lambda}^{\mathrm{cv}}=\frac{1}{n}\sum_{i=1}^{n}W^{\mathrm{cv}}_{i}+o\left(\frac{1}{\sqrt{n\log n}}\right).$
Similarly, there exist i.i.d. Gaussian random variables $(W^{\mathrm{cv+}}_{i})_{i\geq 1}$ with mean $\gamma$ and variance
$\nu^{\mathrm{cv+}}:=\nu_{\lambda^{*}}^{\mathrm{cv}}=\mathrm{var}(V)\left[1-(1-r)\rho_{U,V}^{2}\right]$
such that, almost surely as $n\to\infty$,
$\widehat{\gamma}^{\mathrm{cv+}}=\frac{1}{n}\sum_{i=1}^{n}W^{\mathrm{cv+}}_{i}+o\left(\frac{1}{\sqrt{n\log n}}\right).$
The estimators $\widehat{\nu}_{\lambda}^{\mathrm{cv}}((U_{i},V_{i})_{i=1}^{n},(\widetilde{U}_{i})_{i=1}^{N_{n}})$ and $\widehat{\nu}^{\mathrm{cv+}}((U_{i},V_{i})_{i=1}^{n})$ are consistent estimators of $\nu^{\mathrm{cv}}$ and $\nu^{\mathrm{cv+}}$, respectively.
Proof.We have
@@
The random variables $\overline{V}-\lambda(\overline{U}-\mu)$ and $\lambda(\widehat{\mu}-\mu)$ are independent and are both sample average of i.i.d. random variables with finite moment of order $q$ for some $q>2$. By KMT strong coupling (Theorem S4), there exist i.i.d. Gaussian random variables $(G^{(1)}_{i})_{i\geq 1}$ with mean $\gamma$ and variance $\mathrm{var}(V-\lambda U)$ such that
@@
If $r>0$, then, by Proposition S1, there exist i.i.d. Gaussian random variables $(G^{(2)}_{i})_{i\geq 1}$, independent of $(G^{(1)}_{i})_{i\geq 1}$, with mean 00 and variance $r\lambda^{2}\mathrm{var}(U)$ such that,
@@
Setting $W^{\mathrm{cv}}_{i}=G^{(1)}_{i}$ if $r=0$ and $W^{\mathrm{cv}}_{i}=G^{(1)}_{i}+G^{(2)}_{i}$ if $r>0$ gives the first result, as $\frac{1}{t^{1-1/(2+\delta)}}=o\left(\frac{1}{\sqrt{t\log t}}\right)$. From this, we deduce the second result with Proposition S4, noting that
$\nu^{\mathrm{cv+}}:=\nu_{\lambda^{*}}^{\mathrm{cv}}=\mathrm{var}(V)\left[1-(1-r)\rho_{U,V}^{2}\right].$ (S46)
We prove the consistency of the variance estimators exactly as in Proposition 4.2. ∎
S4.4 Tighter results on prediction-powered inference
Consider the notations of Section 5. Assume that, for any $\theta\in\mathbb{R}$ and $i\geq 1$, we have $\mathbb{E}|U_{\theta,i}|^{2+\delta},\mathbb{E}|V_{\theta,i}|^{2+\delta}<\infty$ for some $\delta>0$, and that $|\frac{n}{N_{n}}-r|=O(1/n^{1-a})$ with $0<a<2/(2+\delta)$ for some $r\in[0,1]$. Then, the AsympCS defined in Proposition 5.1, Proposition 5.2, Proposition 5.3 and Equation 22 in Section 5 have approximation rate $1/\sqrt{n\log{n}}$. This follows trivially from an application of Theorem S2, Theorem S3 and Proposition S5 to the estimators involved, which are either sample mean or control-variate estimators.
@@
$\Pr(\mu\in\mathcal{C}_{\alpha,t}\text{ \leavevmode\nobreak\ for all }t\geq 1)\geq 1-\alpha.$
We now introduce the notion of an asymptotic spherical confidence sequence, inspired by (Waudby-Smith et al., 2024b, Section B.10).
@@
(Asymptotic Spherical Confidence Sequence) Let $\alpha\in(0,1)$. Let $(\widehat{\mu}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\mu$. The sequence of random balls $(\mathcal{C}_{\alpha,t})_{t\geq 1}$, with $\mathcal{C}_{\alpha,t}=B(\widehat{\mu}_{t},R_{t})$ and $R_{t}>0$, is said to be an asymptotic spherical confidence sequence, if there exists a (usually unknown) confidence sequence $(\mathcal{C}^{\star}_{\alpha,t})_{t\geq 1}$, with $\mathcal{C}^{\star}_{\alpha,t}=B(\widehat{\mu}_{t},R_{t}^{\star})$, such that
@@
$R_{t}/R^{\star}_{t}\to 1\text{ a.s. as }t\to\infty.$
We say than an AsympCS has approximation rate $a_{t}$, with $\lim_{t\to\infty}a_{t}=0$, if $|R_{t}-R^{\star}_{t}|=O(a_{t})\text{\leavevmode\nobreak\ a.s. as\leavevmode\nobreak\ }t\rightarrow+\infty.$
S5.2 Nonasymptotic Bayes-assisted CS for i.i.d. Gaussian random vectors
[…]
$\mathcal{C}_{\alpha,t}(\overline{Y}_{t},\Sigma;\pi)=\left\{\mu\in\mathbb{R}^{p}\mid\|\Sigma^{-1/2}(\mu-\overline{Y}_{t})\|\leq\frac{1}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}\eta_{t}(\Sigma^{-1/2}\overline{Y}_{t})^{2}}\right)}\right\}$
forms a $(1-\alpha)$ confidence sequence for $\mu$. One could also consider spherical confidence intervals using $\Lambda_{\max}(\Sigma)$, the maximum eigenvalue of $\Sigma$, similarly to what is done in (Waudby-Smith et al., 2024b, Section B10) for non-assisted confidence regions. In this case, the corresponding (more conservative) $(1-\alpha)$ confidence sequence for $\mu$ is
$\mathcal{C}^{\mathrm{\mathtt{m}BA}}_{\alpha,t}(\overline{Y}_{t},\Sigma;\pi)=\left\{\mu\in\mathbb{R}^{p}\mid\|\mu-\overline{Y}_{t}\|\leq\frac{\sqrt{\Lambda_{\max}(\Sigma)}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}\eta_{t}(\Sigma^{-1/2}\overline{Y}_{t})^{2}}\right)}\right\}.$ (S47)
S5.3 Multivariate extension of Theorems 3.5 and 3.6
To illustrate that our results extend to the multivariate case, we provide multivariate (and tight) versions of Theorem 3.6 and Theorem 3.5.
@@
(Multivariate version of Theorem 3.6) Let $(Y_{t})_{t\geq 1}$ be i.i.d. random vectors in $\mathbb{R}^{d}$ with mean $\mu$ and covariance matrix $\Sigma$. Let $(\widehat{\mu}_{t})_{t\geq 1}$ be a sequence of consistent estimators of $\mu$. Assume that there exists a sequence of i.i.d. Gaussian vectors with mean $\mu$ and covariance matrix $\Sigma$ such that, a.s. as $t\rightarrow\infty$,
$\widehat{\mu}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+\varepsilon_{t}\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ where\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\varepsilon_{t}=o\left(\frac{1}{\sqrt{t\log t}}\right).$ (S48)
Let $(\widehat{\Sigma}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\Sigma$ and $\pi$ be a continuous and proper prior density on $\mathbb{R}^{d}$. Then, $\mathcal{C}^{\mathrm{\mathtt{m}BA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\Sigma}_{t};\pi)$ forms a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$.
Proof.(Theorem S3) By hypothesis, we have, almost surely,

### v2

@@ 7 Discussion
$U_{\alpha,t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}-\varepsilon_{t},\text{~and}$
$L_{\alpha,t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\left(1+\frac{1}{t\rho^{2}}\right)\log\left(\frac{t\rho^{2}+1}{\alpha^{2}}\right)}+\varepsilon_{t},$
for the non-assisted case, and
$U_{\alpha,t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log{\eta}_{t}\left(\frac{\widehat{\mu}_{t}-\varepsilon_{t}}{\sigma}\right)}-\varepsilon_{t},\text{~and}$
@@
$L_{\alpha,t}^{\ast}$ $=\frac{\sigma}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}}\right)-2\log{\eta}_{t}\left(\frac{\widehat{\mu}_{t}-\varepsilon_{t}}{\sigma}\right)}+\varepsilon_{t},$
for the Bayes-assisted case, where $\varepsilon_{t}=\widehat{\mu}_{t}-\overline{W}_{t}=o(a_{t})$ a.s. does not depend on $\alpha$. In both cases, $(\mathcal{C}^{\star}_{\alpha,t})_{t\geq 1}$ is an exact CS (Theorems˜S10 and S11); hence, for $E_{\alpha}^{\star}=\{\mu\in\mathcal{C}^{\star}_{\alpha,t}\text{ for all }t\geq 1\}$,
$\Pr(E_{\alpha}^{\star})\geq 1-\alpha.$
Additionally,
$U_{\alpha,t}\sim U^{\star}_{\alpha,t}\sim L^{\star}_{\alpha,t}\sim\sigma\sqrt{\frac{\log t}{t}}\text{ a.s. as }t\to\infty,$
and, as shown in the proofs of Theorems˜2 and 4, a.s.,
$U_{\alpha,t}-U^{\star}_{\alpha,t}=o(a_{t})\text{ and }U_{\alpha,t}-L^{\star}_{\alpha,t}=o(a_{t}).$ (S38)
Let $\alpha^{\prime}\in(\alpha,1)$. We now aim to show that, for some random, finite time $T_{\alpha^{\prime}}$, $\mathcal{C}^{\star}_{\alpha^{\prime},t}\subseteq\mathcal{C}_{\alpha,t}$ for all $t\geq T_{\alpha^{\prime}}$. We have
$U^{\star}_{\alpha,t}-U^{\star}_{\alpha^{\prime},t}=\frac{(U^{\star}_{\alpha,t})^{2}-(U^{\star}_{\alpha^{\prime},t})^{2}}{U^{\star}_{\alpha,t}+U^{\star}_{\alpha^{\prime},t}}\sim\frac{\sigma^{2}/t\log\frac{(\alpha^{\prime})^{2}}{\alpha^{2}}}{2\sigma\sqrt{\log t/t}}\sim c(\alpha,\alpha^{\prime})a_{t}$ (S39)
a.s. as $t\to\infty$, where $c(\alpha,\alpha^{\prime})=\sigma\log(\alpha^{\prime}/\alpha)>0$. Similarly,
$L^{\star}_{\alpha,t}-L^{\star}_{\alpha^{\prime},t}\sim c(\alpha,\alpha^{\prime})a_{t}\text{ a.s. as }t\to\infty.$ (S40)
Combining Equations˜S40 and S39 with Equation˜S38, we obtain, a.s.
$U_{\alpha,t}-U^{\star}_{\alpha^{\prime},t}\sim c(\alpha,\alpha^{\prime})a_{t}$ (S41)
$L_{\alpha,t}-L^{\star}_{\alpha^{\prime},t}\sim c(\alpha,\alpha^{\prime})a_{t}.$ (S42)
So, there exists a collection of events $\Omega_{0}$ with $\Pr(\Omega_{0})=1$ such that for every $\omega\in\Omega_{0}$, there is a finite $T_{\alpha^{\prime}}(\omega)$ with
$U_{\alpha,t}(\omega)-U^{\star}_{\alpha^{\prime},t}(\omega)\geq\frac{1}{2}c(\alpha,\alpha^{\prime})a_{t}$ (S43)
$L_{\alpha,t}(\omega)-L^{\star}_{\alpha^{\prime},t}(\omega)\geq\frac{1}{2}c(\alpha,\alpha^{\prime})a_{t}.$ (S44)
for all $t\geq T_{\alpha^{\prime}}(\omega)$. Therefore, $\mathcal{C}^{\star}_{\alpha^{\prime},t}\subseteq\mathcal{C}_{\alpha,t}$ for all $t\geq T_{\alpha^{\prime}}$. It follows, that for every $m\geq 1$,
$\Pr(\mu\in\mathcal{C}_{\alpha,t}\text{ for all }t\geq m)\geq\Pr(E^{\star}_{\alpha^{\prime}}\cap\{T_{\alpha^{\prime}}\leq m\})\longrightarrow_{m\to\infty}\Pr(E^{\star}_{\alpha^{\prime}})\geq 1-\alpha^{\prime}.$
Hence, for any $\alpha^{\prime}\in(\alpha,1)$, $\lim\inf_{m\to\infty}\Pr(\mu\in\mathcal{C}_{\alpha,t}\text{ for all }t\geq m)\geq 1-\alpha^{\prime}$ thus
${\lim\inf}_{m\to\infty}\Pr(\mu\in\mathcal{C}_{\alpha,t}\text{ for all }t\geq m)\geq 1-\alpha.$
S3.4 Proof of Proposition˜1
Let
$\epsilon_{n}$ $=\widehat{\gamma}^{\mathrm{cv+}}-\widehat{\gamma}_{\lambda^{\star}}^{\mathrm{cv}}$
$)\displaystyle=\widehat{\gamma}^{\mathrm{cv+}}-(\overline{V}-\lambda^{\star}(\overline{U}-\widehat{\mu}$))
$=(\lambda^{\star}-\widehat{\lambda})(\overline{U}-\mu)-(\lambda^{\star}-\widehat{\lambda})(\widehat{\mu}-\mu).$
@@
$|\epsilon_{n}|\leq|\lambda^{\star}-\widehat{\lambda}|(|\overline{U}-\mu|+|\widehat{\mu}-\mu|).$
By the Lemma˜S1 we have
$|\widehat{\lambda}-\lambda^{\star}|=o\left(n^{-2/(2+\delta)}\right)\text{ a.s. as }n\to\infty$
@@
$|\overline{U}-\mu|=O\left(\sqrt{\frac{\log\log n}{n}}\right),~~~|\widehat{\mu}-\mu|=O\left(\sqrt{\frac{\log\log n}{n}}\right)~~\text{ a.s. as }n\to\infty.$ (S45)
@@
$|\epsilon_{n}|=O\left(\frac{1}{n^{2/(2+\delta)}}\sqrt{\frac{\log\log n}{n}}\right)=o\left(\frac{1}{\sqrt{n\log n}}\right)~~\text{ a.s. as }n\to\infty.$ (S46)
S3.5 Proof of Proposition˜2
We have
@@
The random variables $\overline{V}-\lambda(\overline{U}-\mu)$ and $\lambda(\widehat{\mu}-\mu)$ are independent and are both sample average of i.i.d. random variables with finite moment of order $q$ for some $q=2+\delta>2$. Note that 11/q>1/21-1/q>1/2. By KMT strong coupling (Theorem˜S8), there exist i.i.d. Gaussian random variables $(G^{(1)}_{i})_{i\geq 1}$ with mean $\gamma$ and variance $\mathrm{var}(V-\lambda U)$ such that
@@
If $r>0$, then, by Proposition˜S9, there exist i.i.d. Gaussian random variables $(G^{(2)}_{i})_{i\geq 1}$, independent of $(G^{(1)}_{i})_{i\geq 1}$, with mean 0 and variance $r\lambda^{2}\mathrm{var}(U)$ such that,
@@
Setting $W^{\mathrm{cv}}_{i}=G^{(1)}_{i}$ if $r=0$ and $W^{\mathrm{cv}}_{i}=G^{(1)}_{i}+G^{(2)}_{i}$ if $r>0$ gives the Gaussian coupling (15) for $\nu_{\lambda}^{\mathrm{cv}}$, as $\frac{1}{n^{1-1/q}}=o\left(\frac{1}{\sqrt{n\log n}}\right)$. From this, using Equation˜14, we deduce the coupling (16) for $\widehat{\gamma}^{\mathrm{cv+}}$, noting that
$\nu^{\mathrm{cv+}}:=\nu_{\lambda^{\star}}^{\mathrm{cv}}=\mathrm{var}(V)\left[1-(1-r)\rho_{U,V}^{2}\right].$ (S48)
We have
$\frac{1}{n-2}\sum_{i=1}^{n}(V_{i}-\overline{V}-\lambda(U_{i}-\overline{U}))^{2}\to\mathrm{var}(V-\lambda U)\text{ a.s.}$
and
$\frac{n\lambda^{2}}{N_{n}(N_{n}-1)}\sum_{j=1}^{N_{n}}(\widetilde{U}_{j}-\widehat{\mu})^{2}\to r\lambda^{2}\mathrm{var}(U)\text{ a.s.}$
therefore $\widehat{\nu}_{\lambda}^{\mathrm{cv}}\to\nu_{\lambda}^{\mathrm{cv}}$ a.s. Finally, we show that $\widehat{\nu}^{\mathrm{cv+}}$ is a consistent estimator of $\nu^{\mathrm{cv+}}$. Set $\delta_{i}=V_{i}-\overline{V}-\widehat{\lambda}(U_{i}-\overline{U})$. We have $\delta_{i}=V_{i}-\widehat{\alpha}-\widehat{\beta}U_{i}$ where $\widehat{\alpha}=\overline{V}-\widehat{\lambda}\overline{U}$ and $\widehat{\beta}=\widehat{\lambda}$ are the least squares estimates, minimising $\sum_{i=1}^{n}(V_{i}-\alpha-\beta U_{i})^{2}$. It is well known [21, Theorem 2] that, for
$(\alpha^{\star},\beta^{\star})$ $=\arg\min_{\alpha,\beta}\mathbb{E}[(V-\alpha-\beta U)^{2}]=\left(\gamma-\mu\frac{\mathrm{cov}(U,V)}{\mathrm{var}(U)},\frac{\mathrm{cov}(U,V)}{\mathrm{var}(U)}\right)$
we have
$\frac{1}{n-2}\sum_{i=1}^{n}\delta_{i}^{2}\to\mathbb{E}[(V-\alpha^{\star}-\beta^{\star}U)^{2}]=\mathrm{var}(V)(1-\rho_{U,V}^{2})~~\text{ a.s. as }n\to\infty.$ (S49)
Additionally, by the strong law of large numbers,
$\frac{n/N_{n}}{n-1}\sum_{i=1}^{n}(V_{i}-\overline{V})^{2}\to r\mathrm{var}(V)~~\text{ a.s. as }n\to\infty.$ (S50)
Hence,
almost surely as $n\to\infty$.
S3.6 Proof of Proposition˜4
We apply Theorem˜1 to the i.i.d. sequence $(\ell_{\theta}^{\prime}(\widetilde{X}_{i},f(\widetilde{X}_{i})))_{i\geq 1}$, to obtain an AsympCS $(\widetilde{\mathcal{R}}_{\delta,\theta,i})_{i\geq 1}$ for $m_{\theta}$ with approximation rate $1/\sqrt{i\log{i}}$. The subsequence $(\mathcal{R}_{\delta,\theta,n})_{n\geq 1}$ with $\mathcal{R}_{\delta,\theta,n}=\widetilde{\mathcal{R}}_{\delta,\theta,N_{n}}$ is also an AsympCS for $m_{\theta}$ with approximation rate $1/\sqrt{n\log{n}}$. Asymptotic Type-I error control follows directly from Theorem˜5.
S3.7 Proof of Proposition˜5
In the PPI case, the proof follows from a direct application of Theorem˜1 (non-assisted) or Theorem˜3 (Bayes-assisted) to the sequence of i.i.d. random variables $(V_{\theta,i}-U_{\theta,i})_{i=1}^{n}$. In the PPI++ case, it follows from an application of Theorem˜2 (non-assisted) or Theorem˜4 (Bayes-assisted), together with Proposition˜2, to the control variate estimator (23). Asymptotic Type-I error control follows directly from Theorem˜5.
@@
$\Pr(\mu\in\mathcal{C}_{\alpha,t}\text{ ~for all }t\geq 1)\geq 1-\alpha.$
We now introduce the notion of an asymptotic spherical confidence sequence, inspired by [15, Section B.10].
@@
(Asymptotic Spherical Confidence Sequence) Let $\alpha\in(0,1)$ and $(a_{t})_{t\geq 0}$ a real sequence such as $\lim_{t\to\infty}a_{t}=0$. Let $(\widehat{\mu}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\mu$. The sequence of random balls $(\mathcal{C}_{\alpha,t})_{t\geq 1}$, with $\mathcal{C}_{\alpha,t}=B(\widehat{\mu}_{t},R_{t})$ and $R_{t}>0$, is said to be an asymptotic spherical confidence sequence with (little-o) approximation rate $a_{t}$, if there exists a (usually unknown) confidence sequence $(\mathcal{C}^{\star}_{\alpha,t})_{t\geq 1}$, with $\mathcal{C}^{\star}_{\alpha,t}=B(\widehat{\mu}_{t},R_{t}^{\star})$, such that
@@
$|R_{t}-R^{\star}_{t}|=o(a_{t})\text{~a.s. as~}t\rightarrow+\infty.$
S4.2 Nonasymptotic Bayes-assisted CS for i.i.d. Gaussian random vectors
Let $Y_{1},Y_{2},\ldots,$ be i.i.d. Gaussian random vectors with mean $\mu\in\mathbb{R}^{d}$ and known $d$-by-$d$ positive definite covariance matrix $\Sigma$. Let $\uppi$ be some prior on $\Sigma^{-1/2}\mu$ and define $\eta_{t}(z)=\int\mathcal{N}(z;\zeta,I_{d}/t)\uppi(\zeta)d\zeta$ where $I_{d}$ denotes the $d$-by-$d$ identity matrix. By the method of mixtures and Ville’s inequality (similarly as in Theorem˜S11), the sequence of ellipsoid regions defined by
$\mathcal{C}_{\alpha,t}(\overline{Y}_{t},\Sigma;\pi)=\left\{\mu\in\mathbb{R}^{d}\mid\|\Sigma^{-1/2}(\mu-\overline{Y}_{t})\|\leq\frac{1}{\sqrt{t}}\sqrt{\log\left(\frac{t^{d}}{(2\pi)^{d}\alpha^{2}\eta_{t}(\Sigma^{-1/2}\overline{Y}_{t})^{2}}\right)}\right\}$
forms a $(1-\alpha)$ confidence sequence for $\mu$. One could also consider spherical confidence intervals using $\Lambda_{\max}(\Sigma)$, the maximum eigenvalue of $\Sigma$, similarly to what is done in [15, Section B10] for non-assisted confidence regions. In this case, the corresponding (more conservative) $(1-\alpha)$ confidence sequence for $\mu$ is
$\mathcal{C}^{\mathrm{\tt mBA}}_{\alpha,t}(\overline{Y}_{t},\Sigma;\pi)=\left\{\mu\in\mathbb{R}^{d}\mid\|\mu-\overline{Y}_{t}\|\leq\frac{\sqrt{\Lambda_{\max}(\Sigma)}}{\sqrt{t}}\sqrt{\log\left(\frac{t^{d}}{(2\pi)^{d}\alpha^{2}\eta_{t}(\Sigma^{-1/2}\overline{Y}_{t})^{2}}\right)}\right\}.$ (S51)
S4.3 Multivariate extension of Theorems˜3 and 4
To illustrate that our results extend to the multivariate case, we provide multivariate (and tight) versions of Theorem˜4 and Theorem˜3.
@@
(Multivariate version of Theorem˜4) Let $(\widehat{\mu}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\mu$. Assume that there exists a sequence of i.i.d. Gaussian vectors with mean $\mu$ and positive definite covariance matrix $\Sigma$ such that, a.s. as $t\rightarrow\infty$,
$\widehat{\mu}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+\varepsilon_{t}\text{~~~where~~~}\varepsilon_{t}=o\left(\frac{1}{\sqrt{t\log t}}\right).$ (S52)
Let $(\widehat{\Sigma}_{t})_{t\geq 1}$ be a consistent sequence of estimators of $\Sigma$ such that $\|\widehat{\Sigma}_{t}-\Sigma\|=o(1/\log t)$ a.s. where $\|\cdot\|$ is the spectral norm, and $\uppi$ be a continuous and proper prior density on $\mathbb{R}^{d}$. Then, $\mathcal{C}^{\mathrm{\tt mBA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\Sigma}_{t};\uppi)$ forms a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$.
Proof.
(Theorem˜S12) By hypothesis, we have, almost surely,